Disposal of Fixed Asset

Equipment acquired on January 6 at a cost of $358,800, has an estimated useful life of 10 years and an estimated residual value of $46,800.

a. What was the annual amount of depreciation for the Years 1-3 using the straight-line method of depreciation?

Year Depreciation Expense
Year 1 $
Year 2 $
Year 3 $

b. What was the book value of the equipment on January 1 of Year 4?
$

c. Assuming that the equipment was sold on January 3 of Year 4 for $251,900, journalize the entry to record the sale. If an amount box does not require an entry, leave it blank.

d. Assuming that the equipment had been sold on January 3 of Year 4 for $270,500 instead of $251,900, journalize the entry to record the sale. If an amount box does not require an entry, leave it blank.
Answer #1

a. What was the annual amount of depreciation for the Years 1-3 using the straight-line method of depreciation?

Year Depreciation Expense
Year 1 (358800-46800/10) = 31200
Year 2 $31200
Year 3 $31200

b) Book value on January 1, Year 4 = 358800-(31200*3) = 265200

c) Journal entry

Date account and explanation debit credit
jan 3, Year 4 Cash 251900
Accumulated depreciation-Equipment 93600
Loss on sale of equipment 13300
Equipment 358800
(To record sale of equipment)

d) Journal entry

Date account and explanation debit credit
jan 3, Year 4 Cash 270500
Accumulated depreciation-Equipment 93600
Gain on sale of equipment 5300
Equipment 358800
(To record sale of equipment)